# Arrow-Debreu Meets Kyle: Price Discovery for Derivatives

### v26

@@ 7.1 Informed Demand
  • Step 1. Form the initial portfolio by buying $1-\frac{1}{K}$ shares of the observed distribution $\eta(\,\cdot\,|s_{l})$ and selling $\frac{1}{K}$ shares of each remaining distribution $\eta(\,\cdot\,|s_{l})$, $l\neq k$.
  • Step 2. Adjust the initial portfolio allocations by applying the linear transformation ${\bf L}^{-1}$.
  • Step 3. Scale the overall portfolio by the endogenous constant $\alpha^{*}$.
@@
  • Information Intensity ($\frac{\sigma_{\varepsilon}}{\sigma_{v}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation of asset payoff from the variation of order flow, therefore reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can more effectively infer the variation of payoff distributions across signals from the variation of order flow across markets, which reduces informed demand.
@@
The following examples illustrate the informed demand for trading on the first, second, and third moments of an underlying asset payoff. These examples explain how the corresponding option strategies used in trading practice emerge as equilibrium strategies. Other common option strategies are explained similarly under appropriate specifications. In Examples 7.1, 7.2, and 7.3, signal is binary, i.e., $S=\{s_{1},s_{2}\}$, and all payoff distributions are suitably discretized.
@@
($x$-axis: state/strike, $y$-axis: probability)
@@
($x$-axis: state/strike, $y$-axis: payoff)
@@
Bull/Bear Spread The left column shows the insider’s private signal, as indicated by the shaded distributions. The right column shows the informed demand $W^{*}$ conditional on the corresponding signals.
@@
Let signals $s_{1}$ and $s_{2}$ specify normal payoff distributions with respective means $\mu_{1}>\mu_{2}$ and an identical variance. Observing $s_{1}$ (resp. $s_{2}$) informs the insider of a high (resp. low) expected payoff. This is a Kyle (1985)-type setting where private information concerns the expected payoff, but the insider here can trade derivatives.
The left column of Figure 3 shows the possible payoff distributions with the insider’s private signal indicated by the shaded distribution. The right column of Figure 3 shows the corresponding informed demand for the indicated signals.
These nonlinear informed demands require options. Indeed, they correspond to well-known option strategies (see Hull (2015)): Conditional on high expected payoff, the informed demand in Figure 3b takes the form of a bull spread. Conditional on low expected payoff, the informed demand in Figure 3d takes the form of a bear spread. These option strategies maximize gains in the anticipated direction (up or down) while limiting exposure at extreme out-of-the-money strikes that are unlikely to pay off. The linear demand in Kyle (1985) is sub-optimal because it entails unnecessarily large bets at the extreme strikes.
@@
Straddle/Butterfly The left column shows the insider’s private signal, as indicated by the shaded distributions. The right column shows the informed demand $W^{*}$ conditional on the corresponding signals.
Price Impact Three states are indicated—$x_{1}$, $x_{2}$, and $x_{3}$. Security $x_{2}$ has zero price impact on all securities, because its payoff is constant—i.e. has zero variation—across signals. Therefore, the informed demand for $x_{2}$ must be zero, i.e., $W^{*}(x_{2}|s_{1})=W^{*}(x_{2}|s_{2})=0$, as shown by Figures 4b and 7.2 of the right column. Securities $x_{1}$ and $x_{3}$ have positive cross price impact because they have positively correlated—in fact, identical—payoffs.
@@
The left column of Figure 4 shows the possible payoff distributions, with the insider’s private signal indicated by the shaded distribution. The right column of Figure 4 shows the informed demand corresponding to the indicated signals. The informed demands correspond to classical option strategies for exploiting expected volatility (see Hull (2015)): Conditional on high volatility, the informed demand in Figure 4b resembles a straddle or long iron condor pattern. Conditional on low volatility, the informed demand in Figure  follows a butterfly pattern.
@@
Figure 5: Trading on Skewness The left column shows the insider’s private signal, as indicated by the shaded distributions. The right column shows the informed demand $W^{*}$ conditional on the corresponding signals.
@@
Let signals $s_{1}$ and $s_{2}$ specify skew normal payoff distributions with respective skewness parameters $\alpha_{1}>\alpha_{2}$. All other moments besides skewness are equal across signals. The left column of Figure 5 shows the possible payoff distributions with the insider’s private signal indicated by the shaded distribution. The right column of Figure 5 shows the informed demand corresponding to the indicated signals. [^19] In practitioner parlance, the informed demand is implemented by long-short option strategies called “Christmas tree spreads” (see Vine (2011)). The long part of the portfolio is aligned with the direction of skewness.
@@ 7.2 Price Impact
where ${\operatorname{\mathbb{Cov}}}(\,\cdots\,,\,\cdots\,\big{|}\omega)$ is the covariance over signals $s_{k}\in S$ under the market maker’s equilibrium posterior, conditional on order flow $\omega$. The expection $\mathbb{E}[\,\cdot\,]$ is taken over $\omega$ with respect to its equilibrium distribution. Dropping this expectation in (34) gives the equilibrium price impact conditional on $\omega$.
@@
The equilibrium cross price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$ is the covariance of their payoffs, with adjustments for the sizes of informed and noise trades—(i) ,(ii), and (iii) below.
@@
    Payoff Covariance $(\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right])$: The (expected) covariance of the payoffs of securities $x_{i}$ and $x_{j}$ under the market maker’s equilibrium posterior.
Therefore, $\Lambda_{i,j}$ generalizes the intuition of the cross price impact discussed in Section 2, showing this simple intuition of cross-market dynamics to be fully robust.
@@
    (Negative Cross Price Impact.) If security $x_{i}$ pays off only under a signal $s_{k_{i}}$ and security $x_{j}$ only under $s_{k_{j}}\neq s_{k_{i}}$, their equilibrium cross price impact is negative. This is the same scenario as Example 5.3(ii), but in equilibrium. If the market maker receives a buy order for $x_{i}$, he infers that $x_{j}$ is unlikely to pay off and lowers its price accordingly, and vice versa. Indeed,
@@
    (Zero Cross Price Impact.) If a security $x_{j}$ has zero variation in its payoff across signals (i.e., $\eta(x_{j}|\,\cdot\,)$ is constant), then its order flow has zero price impact on all markets. In turn, the informed demand for $x_{j}$ must be zero—by no-arbitrage. Conversely, if the informed demand for $x_{j}$ is zero across signals, its order flow has zero information content and, therefore, zero price impact on all markets. This is exemplified by the security $x_{2}$ indicated in Figure 4c. Figures 4b and 7.2 show the corresponding informed demand $W^{*}(x_{2}|\,\cdot\,)=0$ to be indeed zero across signals.
@@
    (Positive Cross Price Impact.) If two securities have identical payoffs, then, of course, their cross price impact is positive. This is the case for the securities $x_{1}$ and $x_{3}$ indicated in Figure 4c.
We have the following general formula for the cross price impact between derivatives.
@@
Applying Corollary 7.5 to Example 7.2 gives the following result.
@@
In other words, a high cross price impact between the put-call pair in a long-volatility straddle means that volatility is underpriced and should, therefore, correct upwards as volatility price discovery takes place.
@@ 7.3 Information Efficiency of Prices
The probability distribution of the market maker’s equilibrium posterior over order flow $\omega$ has the following characterization.
@@
The equilibrium distribution of AD prices is an immediate corollary (we can similarly characterize the equilibrium $\omega$-by-$\omega$ pricing kernel).
@@
When the insider observes $s_{k}$, the equilibrium AD prices $P^{*}(\,\cdot\,|\omega)$ are distributed as $P^{*}(\,\cdot\,|\omega)\stackrel{{\scriptstyle d}}{{\sim}}\sum\limits_{l=1}^{K}q_{l}^{(k)}\eta(\,\cdot\,|s_{l})$, with $q^{(k)}$ as defined in Proposition 7.7. Therefore, the expected equilibrium AD prices $\overline{P}^{*}(\,\cdot\,)$ are given by
@@
In the Kyle (1985) setting, information efficiency pertains to a single-asset. In contrast, in our setting, information efficiency extends to multiple derivative securities, capturing not only first-moment information but also higher moments. In Kyle (1985), information efficiency is measured by the prior-to-posterior variances ratio of the asset price. This single-asset measure is inadequate in our setting—for example, it is unclear how security-by-security variance ratios would reflect how much volatility information is incorporated across prices. For a single asset, a higher variance ratio means that a larger fraction of the insider’s information is incorporated into the market price. In our setting, information efficiency is summarized directly by the fraction of the insider’s information (i.e., true payoffs) in the market prices.
@@
The information efficiency $\mathbb{E}[q_{k}^{(k)}]$ of AD prices is independent of the possible distributions across future states $\eta(\,\cdot\,|\,\cdot\,)$ and the noise trading intensity $(\sigma_{i})$.
@@
Corollary 7.10 shows a basic property of complete markets that has been unnoticed in prior literature—the information efficiency of prices is robust to the specifics of information. As the insider’s trades (co-)vary across securities according to his information intensity, the market maker correctly anticipates these (co-)variations and equilibrates prices to incorporate the insider’s information. The degree to which this occurs is independent of the specific information held by the insider. This is the price discovery counterpart of efficient risk sharing in complete markets, where prices equilibrate to allocate risk efficiently independent of the initial distribution of risk.
Figure 6 shows the (expected) posterior probabilities $\mathbb{E}[q_{k}^{(1)}]$, $1\!\leq\!k\!\leq\!K$, conditional on the signal $s_{1}$ as $K$ increases. It shows that the information efficiency $\mathbb{E}[q_{1}^{(1)}]=\mathbb{E}[q_{k}^{(k)}]$ is decreasing in the number of signals. More precisely, Figure 7 shows that $\mathbb{E}[q_{k}^{(k)}]$ is a decreasing, convex function of $K$. This is Proposition 7.12.
@@
Proposition 7.12 is intuitive (and consistent with the comparative statics of $\alpha^{*}$). As the market maker’s prior becomes more dispersed over possible signals, he is less able to infer information from trades, which lowers the information efficiency of prices. The marginal effect is diminishing because the larger trades the insider makes to exploit the additional prior uncertainty of the market maker also increases the variation of his trades across signals, which reveals some of his information to partially offset the information efficiency loss.
@@ 8 Empirical Implications
Empirical studies on adverse selection traditionally focus on the within-market price impact, which estimates Kyle’s lambda. [^20] See Glosten and Harris (1988), Lin et al. (1995), Huang and Stoll (1997), Goyenko et al. (2009), Hendershott et al. (2011), and Makarov and Schoar (2020). However, this approach does not capture information regarding higher moments. For volatility information, the implied volatility from option prices serves as a common proxy for future volatility expectations. [^21] See, for example, Goyal and Saretto (2009). However, implied volatility is based on the assumption of constant volatility and is a result of reduced form model mis-specification. Our characterization of cross price impact extends Kyle’s lambda to an empirical measure that directly reflects higher-moment adverse selection effects across options. This new measure extends existing empirical proxies and can help reduce biases associated with ignoring cross-market effects.
@@
Figure 8: Insider-Induced Volatility Smile Figure 8a shows the strike structure of the implied volatility of the equilibrium option prices in the setting of Example 7.2. For an order flow realization $\omega$, we apply the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to calculate the option prices across strike $K$. Based on these option prices, we calculate the implied volatilities $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plot displaying the average implied volatilities over log-moneyness. (The risk-free rate and dividend yield are set to zero.) Figure 8b shows the insider-induced volatility smiles for $\sigma_{1}$ values of $0.25$ (solid line, as in Figure 8a), $0.3$ (dashed line), and 0.35 (dotted line).

### v27

@@ 7.1 Informed Demand
  • Step 1: Form the initial portfolio by buying $1-\frac{1}{K}$ shares of the observed distribution $\eta(\,\cdot\,|s_{l})$ and selling $\frac{1}{K}$ shares of each remaining distribution $\eta(\,\cdot\,|s_{l})$, $l\neq k$.
  • Step 2: Adjust the initial portfolio allocations by applying the linear transformation ${\bf L}^{-1}$.
  • Step 3: Scale the portfolio by the endogenous constant $\alpha^{*}$.
@@
  • Information Intensity ($\frac{\sigma_{\varepsilon}}{\sigma_{v}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation in the asset payoff from the variation in order flow, thus reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can better infer the variations in the payoff distributions across signals from the variations in cross-market order flow, which reduces informed demand.
@@
The following examples illustrate the informed demand for trading on the first, second, and third moments of an underlying asset payoff. These examples explain how the corresponding option strategies used in trading practice emerge as equilibrium strategies. Other common option strategies are explained similarly under appropriate specifications. In Examples 7.1, 7.2, and 7.3, the signal is binary, i.e., $S=\{s_{1},s_{2}\}$, and all payoff distributions are suitably discretized.
@@
($x$-axis: state/option strike, $y$-axis: probability)
@@
($x$-axis: state/option strike, $y$-axis: payoff)
@@
Bull/Bear Spreads The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
(a)/(b): high expected payoff signal/bull spread, (c)/(d): low expected payoff signal/bear spread.
@@
Let signals $s_{1}$ and $s_{2}$ specify normal payoff distributions with respective means $\mu_{1}>\mu_{2}$ and an identical variance. Observing $s_{1}$ (resp. $s_{2}$) informs the insider of a high (resp. low) expected payoff. This is a Kyle (1985)-type setting, where private information concerns the expected payoff, but the insider here can trade derivatives.
The left column of Figure 3 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 3 shows the insider’s demand corresponding to each signal. These nonlinear informed demands require options. Indeed, they correspond to well-known option strategies (see Hull (2015)). When the insider anticipates high expected payoff, his option strategy in Figure 3b takes the form of a bull spread. When he anticipates low expected payoff, his option strategy in Figure 3d takes the form of a bear spread. These option strategies maximize gains in the anticipated direction (up or down) while limiting exposure at extreme out-of-the-money strikes that are unlikely to pay off. In contrast, the linear demand in Kyle (1985) is sub-optimal because it entails unnecessarily large bets at the extreme strikes.
@@
Straddle/Butterfly The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
(a)/(b): the high volatility signal/straddle, (c)/(d): low volatility signal/butterfly.
Price Impact Three states are indicated—$x_{1}$, $x_{2}$, and $x_{3}$.
$\bullet\,x_{2}$: zero price impact on all securities because its payoff is constant (has zero variation) across signals. Thus, the informed demand for $x_{2}$ must be zero—see (b) and (d)
$\bullet\,x_{1}$ and $x_{3}$: positive cross price impact because they have positively correlated (in fact, identical) payoffs.
@@
The left column of Figure 4 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 4 shows the informed demand corresponding to each signal. The informed demands align with common option strategies for volatility trading (see Hull (2015)). When the insider anticipates high volatility, his option strategy in Figure 4b resembles a straddle or long iron condor pattern. When he anticipates low volatility, his option strategy in Figure 4d follows a butterfly pattern.
@@
Figure 5: Trading on Skewness The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
@@
Let signals $s_{1}$ and $s_{2}$ specify skew-normal payoff distributions with respective skewness parameters $\alpha_{1}>\alpha_{2}$. All moments other than skewness are identical across signals. The left column of Figure 5 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 5 shows the informed demand corresponding to each signal. The informed demand is implemented by long-short option strategies, with the portfolio’s long position aligned with the skewness direction. [^19] In practitioner parlance, these option strategies are called “Christmas tree spreads” (see Vine (2011)).
@@ 7.2 Price Impact
where ${\operatorname{\mathbb{Cov}}}(\,\cdots\,,\,\cdots\,\big{|}\omega)$ is the covariance over signals $s_{k}\in S$ under the market maker’s equilibrium posterior, conditional on order flow $\omega$. The expectation $\mathbb{E}[\,\cdot\,]$ is taken over $\omega$ with respect to its equilibrium distribution. Removing this expectation in (34) yields the equilibrium price impact conditional on $\omega$.
@@
The equilibrium cross price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$ is the covariance of their payoffs, adjusted by factors related to informed and noise trades: (i), (ii), and (iii) below.
@@
    Payoff Covariance $(\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right])$: The (expected) covariance of securities $x_{i}$ and $x_{j}$ payoffs under the market maker’s equilibrium posterior.
Thus, $\Lambda_{i,j}$ extends the intuition of the cross price impact discussed in Section 2, showing the robustness of this basic intuition for cross-market information dynamics.
@@
    (Negative Cross Price Impact.) If security $x_{i}$ pays off only under a signal $s_{k_{i}}$ and $x_{j}$ only under $s_{k_{j}}\neq s_{k_{i}}$, their equilibrium cross price impact is negative. This is the same scenario as Example 5.3(ii), but in equilibrium. Upon receiving a buy order for $x_{i}$, the market maker infers that $x_{j}$ is unlikely to pay off and lowers its price accordingly, and vice versa. Indeed,
@@
    (Zero Cross Price Impact.) If a security $x_{j}$’s payoff has zero variation across signals (i.e., $\eta(x_{j}|\,\cdot\,)$ is constant), then its order flow has zero price impact across markets. In turn, no-arbitrage implies that informed demand for $x_{j}$ must be zero. Conversely, if the informed demand for $x_{j}$ is zero across signals, its order flow has zero information content and, therefore, zero price impact on all markets. This is exemplified by the security $x_{2}$ indicated in Figure 4c. Figures 4b and 4d show the corresponding informed demand $W^{*}(x_{2}|\,\cdot\,)=0$ to be indeed zero across signals.
@@
    (Positive Cross Price Impact.) If two securities have identical payoffs, then their cross price impact is naturally positive. This is the case for the securities $x_{1}$ and $x_{3}$ indicated in Figure 4c.
The general formula for cross price impact between derivatives is as follows.
@@
Applying Corollary 7.5 to Example 7.2 yields the following result.
@@
In other words, a high cross price impact between the put-call pair in a long-volatility straddle means that volatility is underpriced and should therefore correct upwards as volatility price discovery takes place. These insights extend to higher moments beyond volatility, offering testable empirical predictions (see Section 8).
@@ 7.3 Information Efficiency of Prices
The equilibrium posterior distribution of the market maker over order flow $\omega$ is characterized as follows.
@@
The equilibrium distribution of AD prices is an immediate corollary (the equilibrium $\omega$-by-$\omega$ pricing kernel can be characterized similarly):
@@
When the insider observes $s_{k}$, the equilibrium AD prices $P^{*}(\,\cdot\,|\omega)$ are distributed as $P^{*}(\,\cdot\,|\omega)\stackrel{{\scriptstyle d}}{{\sim}}\sum\limits_{l=1}^{K}q_{l}^{(k)}\eta(\,\cdot\,|s_{l})$, with $q^{(k)}$ as defined in Proposition 7.7. Therefore, the expected equilibrium AD prices, $\overline{P}^{*}(\,\cdot\,)$, are given by
@@
In our setting, information efficiency encompasse higher-moment information across multiple securities, unlike Kyle (1985), which considers only the first moment and a single asset. In Kyle (1985), information efficiency is measured by the prior-to-posterior variance ratio of the asset price, where a higher ratio indicates that a larger fraction of the insider’s information is reflected in the price. This single-asset measure is inadequate for our setting. Security-by-security variance ratios do not reflect how much higher-moment information is incorporated across security prices. In our setting, information efficiency is directly measured by the fraction of the insider’s information (i.e., true payoffs) in the market prices.
@@
The information efficiency of AD prices does not depend on the possible distributions across future states $\eta(\,\cdot\,|\,\cdot\,)$ or on the noise trading intensities $(\sigma_{i})$.
@@
Corollary 7.10 points to a basic property of complete markets overlooked in prior literature: the information efficiency of prices is robust to the specifics of information. As the insider’s trades (co-)vary across securities according to his information intensity, the market maker correctly anticipates these (co-)variations and equilibrates prices to incorporate that information. The degree to which this occurs is independent of the specifics of information. This is the price discovery counterpart of efficient risk sharing in complete markets, where prices equilibrate to allocate risk efficiently independent of the initial distribution of risk.
Figure 6 illustrates the expected posterior probabilities $\mathbb{E}[q_{k}^{(1)}]$, $1\!\leq\!k\!\leq\!K$, when the true signal is $s_{1}$, as $K$ increases. It shows that the information efficiency $\mathbb{E}[q_{1}^{(1)}]$ (equal across $k$) is decreasing in the number of signals. More precisely, Figure 7 demonstrates that $\mathbb{E}[q_{k}^{(k)}]$ is a decreasing, convex function of $K$, as stated in Proposition 7.12.
@@
Proposition 7.12 is intuitive and consistent with the comparative statics of $\alpha^{*}$. As the market maker’s prior becomes more dispersed over possible signals, he is less able to infer information from trades, which lowers the information efficiency of prices. The marginal effect is diminishing because the larger trades the insider makes to exploit the additional prior uncertainty of the market maker also increases the variation of his trades across signals. This increased variation reveals some of his information, partially offsetting the loss in information efficiency.
@@ 8 Empirical Implications
Empirical studies on adverse selection traditionally focus on the within-market price impact, which estimates Kyle’s lambda. [^20] See Glosten and Harris (1988), Lin et al. (1995), Huang and Stoll (1997), Goyenko et al. (2009), Hendershott et al. (2011), and Makarov and Schoar (2020). However, this approach does not capture information about higher moments. For volatility information, the implied volatility from option prices serves as a common proxy for future volatility expectations. [^21] See, for example, Goyal and Saretto (2009). However, implied volatility is based on the assumption of constant volatility and is a result of reduced form model mis-specification. Our characterization of cross price impact extends Kyle’s lambda to an empirical measure that directly reflects higher-moment adverse selection effects across options. This new measure extends existing empirical proxies and can help reduce biases associated with ignoring cross-market effects.
@@
Figure 8: Insider-Induced Volatility Smile
(a) displays the insider-induced implied volatility across option strikes under Example 7.2.
(b) displays this insider-induced volatility smiles for different values of true volatility: $0.25$ (solid line, same line as (a)), $0.3$ (dashed line), and $0.35$ (dotted line).
Computation: For an order flow realization $\omega$, we use the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness.
The risk-free rate and dividend yield are set to zero.